An IV solution containing 0.025% of Drug X has been infusing since 6 AM at 100 ml/hr.
How much drug X will the patient receive by noon?
Answer #1

Iv solution containing 0.025% of Drug X has been infusing since 6 AM to noon ( 12 noon) at 100 ml/hr.

Total time = (12-6) hr = 6 hour.

Infusion of drug in 1 hour is 100 ml.

So, in 6 hour it will be (100×6) ml = 600 ml.

0.025 % of drug means 0.025 ml of drug concentration contained in 100 ml infusion.

So , in 600 ml the amount of drug X will be ( 0.025× 6) ml = 0.15 ml.

Therefore the patient receive 0.15 ml of drug by noon.